OpenStudy (1230kent):

2 2/3 * 1 1/3 =10 2/3 is this right

OpenStudy (anonymous):

That is not right.

OpenStudy (anonymous):

The correct answer is 32/9.

OpenStudy (anonymous):

Do you know how to simplify it into a mixed number?

OpenStudy (jdoe0001):

\(\bf 2\cfrac{2}{3}\cdot 1\cfrac{1}{3}\implies \cfrac{2\cdot 3+2}{3}\cdot \cfrac{1\cdot 3+1}{3}\implies \cfrac{(2\cdot 3+2)+(1\cdot 3+1)}{3\cdot 3}\)

OpenStudy (jdoe0001):

hmm actually \(\bf 2\cfrac{2}{3}\cdot 1\cfrac{1}{3}\implies \cfrac{2\cdot 3+2}{3}\cdot \cfrac{1\cdot 3+1}{3}\implies \cfrac{(2\cdot 3+2)(1\cdot 3+1)}{3\cdot 3}\)

OpenStudy (tkhunny):

2 2/3 * 1 1/3 = 2*1 + 2*(1/3) + 1*(2/3) + (2/9) = 2 + (2/3) + (2/3) + (2/9) = 2 + 4/3 + 2/9 = 3 + 1/3 + 2/9 = 3 + 3/9 + 2/9 = 3 5/9 I don't really recommend it as a good method, but it is good practice at manipulating fractions and multiplying binomials.